# UnMarker: A Universal Attack on Defensive Watermarking

### v1

@@ 3. Threat Model
Providers’ Goals. Providers wish to ensure their services are not used to spread disinformation. Thus, they watermark all released images. If suspicious images surface, providers invoke their detectors to test if they contain the watermark, proving them fake.
@@
Providers’ Capabilities. Providers have sufficient resources and data to train SOTA generators. As is common practice (Goo, [n. d.]; DAL, [n. d.]), their APIs allow restricted access, accepting queries (e.g., text) and returning watermarked images without revealing intermediate results.
@@
Attacker’s Goals. Attackers attempt to generate real-looking harmful images by invoking the providers’ APIs and then using methods to 1) remove watermarks while 2) retaining quality.
@@
Attacker’s Capabilities: Resources. Following the defensive watermarking assumptions (see §2), attackers lack high-quality data or extensive resources. Thus, training advanced models or surrogates for the providers’ generators (Lukas et al., 2024) or detectors (Saberi et al., 2024) is not possible. As some services (aws, [n. d.]) allow renting advanced hardware, they may access powerful machines. Yet, this can only be short-term due to high costs, not for training large models over numerous GPU hours.
@@
Black-Box. As many providers keep their algorithms and datasets secret (closed-source) (Saharia et al., 2022), attackers have no whitebox access to (similar) models, data, or any information about the system.
Query-Free. It is unclear how defensive watermarking will be used as commercial systems are still in development or testing (Gowal, 2023). Yet, recent announcements (David and Heath, 2024) suggest detectors may not be exposed publicly but only accessible to providers and chosen entities to identify and purge fake content, forcing attacks to be developed without detector feedback and disqualifying query-based methods (Jiang et al., 2023).
4. Design Philosophy
@@
Since the detector $\mathcal{D}$ must be able to verify the watermark in any image, it should know where to find this watermark, meaning there must be a set of attributes, namely the universal carrier, that can be measured in all images to make this decision. Our analysis will reveal that identifying this set is possible since the requirements from a robust scheme eliminate all options except the spectral amplitudes. Thus, any robust scheme will inject its watermarks in the collective magnitudes of the different frequency bands. We will also show how identifying this carrier enables attackers to mount destructive attacks that remove the watermark with minimal visual impact and no information regarding the system (see §5).
4.1. Uncovering The Universal Carrier
A detector $\mathcal{D}$ is a computationally limited algorithm that can only monitor a finite set of attributes $\mathcal{\boldsymbol{P}}=\{p_{b}|\;b\in[\![B]\!]\}$ (for some $B$) in all images. Deciding if the image is watermarked is based on whether (a subset of) $\mathcal{\boldsymbol{P}}$’s attributes follow a watermark property distribution $\mathcal{\boldsymbol{P}}_{w}$ that $\mathcal{D}$ associates with the watermark. Thus, while $\mathcal{\boldsymbol{P}}$ is a set of attributes, $\mathcal{\boldsymbol{P}}_{w}$ is a distribution over $\mathcal{\boldsymbol{P}}$’s values that identifies the watermark. Practically, there may be several patterns $\mathcal{\boldsymbol{P}}_{w}$ pertaining to different subsets of $\mathcal{\boldsymbol{P}}$. We term $\mathcal{\boldsymbol{P}}$ the universal carrier since it contains all the properties watermarks can influence.
@@ 4.1. Uncovering The Universal Carrier
Although $\mathcal{\boldsymbol{P}}_{w}$ should allow watermarking images with different contents, its variance must still be sufficiently low to prevent detecting $w$ in non-watermarked images, restricting it to a reasonably narrow vicinity. Thus, relatively small deviations in $\mathcal{\boldsymbol{P}}$’s values will cause any watermarked image to violate this distribution $\mathcal{\boldsymbol{P}}_{w}$, removing the watermark without compromising quality. This attack requires no knowledge of the specific $\mathcal{\boldsymbol{P}}_{w}$, making it executable without feedback from $\mathcal{D}$ as it is merely destructive to the channel that carries the watermark. Attackers must identify the carrier $\mathcal{\boldsymbol{P}}$ and then devise a method to disrupt it in a quality-preserving manner, which depends on the content. To identify the carrier $\mathcal{\boldsymbol{P}}$, we first analyze the trivial, widely-accepted requirements (Wen et al., 2023; Lukas and Kerschbaum, 2023; Zhu et al., 2018; Tancik et al., 2020; Yu et al., 2022, 2021; Fernandez et al., 2023) from a robust scheme:
@@
Robust Watermarking Requirements. Universal Applicability: The scheme must be able to watermark images of various concepts and scenes, and the watermark must always be detectable.
@@
Universal Stealthiness. The process cannot introduce unnatural artifacts that degrade image quality. We do not require the watermark not to leave visible traces since semantic watermarks violate this concept by definition, allowing them to influence the structure. Yet, as users can only observe the final, watermarked images, image quality becomes the only metric for quantifying stealthiness.
@@ 4.1.1. Candidates
An image is characterized by its 1) spatial structure—the specific pixel values, and 2) frequency distribution or the rates at which values alternate between neighboring pixels. Thus, the carrier must reside in these channels.
@@
Spatial Structure: Watermarks cannot be embedded in the pixel values as they make for an unreliable carrier that experiences sporadic variations (in the absence of a watermark) between images based on their contents (different images assign extremely different values to their pixels). However, the scheme must always bring the carrier’s values to the same distribution $\mathcal{\boldsymbol{P}}_{w}$ for any watermarked image. When these values naturally (i.e., without the watermark) exhibit a high degree of content dependency, this is impossible without restricting the scheme to processing a small set of similar contents or leaving unnatural cues in many outputs corresponding to contents that would otherwise require these values in images portraying them to lie far away from the desired $\mathcal{\boldsymbol{P}}_{w}$. As this distance increases, the magnitudes of the changes needed to bring them to $\mathcal{\boldsymbol{P}}_{w}$ grow, failing to integrate naturally with the content. Restrictions violate applicability, while unnatural artifacts are not stealthy.
The scheme may strategically place the pixels encoding the watermark to accommodate large modifications with minimal visibility. Two possibilities exist: either they are far apart to remain unnoticeable despite significantly differing from their environments or located in regions the content is unlikely to occupy (i.e., the background). Both are volatile and cannot withstand manipulations (e.g., cropping or noise), violating the robustness requirement (see §6.2.1).
Figure 1. Spectral analysis of two images. Both depict worldly objects of high consistency. Thus, the collective spectral magnitudes are distributed similarly as the low frequencies (center) corresponding to gradual variations are always far more dominant. Phases determine the spatial arrangement of the pixel value shifts that constitute these magnitudes to shape the content, making them extremely different.
Frequency Distribution: There are two options for influencing frequencies to inject watermarks: Locally or globally. Local modifications affect the rates at which pixels vary in select spatial regions. These share the above pitfalls, leaving only global changes to the spectrum. The spectrum associates frequencies with two properties: magnitude and phase. A frequency’s magnitude represents the prevalence of shifts that occur at this rate across the image. Mathematically, any image is a superposition of periodic signals with varying frequencies occurring at different phases that determine their orientations and form the image’s structure based on how they intersect. Phases vary drastically from image to image as the primary factor affecting the structure (see Fig. 1). Thus, they make for an unreliable carrier as well due to this sensitivity to the content.
Amplitudes are collective measurements with little dependency on specific pixels, making them resistant to reasonable manipulations that would preserve the majority of energy (amplitudes) distribution. They are also distributed similarly for any realistic content: lower frequencies contain the majority of energy as they correspond to gradual pixel changes that capture the inherent consistency (continuity) of the worldly objects images depict. Higher frequencies or rapid transitions occur at edges (intersections of multiple objects), making them far less prevalent (see Fig. 1). Thus, while amplitudes vary between images, they always follow a similar distribution, decaying as the frequencies grow. This stability that other alternatives lack allows injecting watermarks via changes of similar magnitudes for any content, enabling them to be minimized to an extent where they can always be integrated naturally.
4.1.2. Spectral Watermarks
The above entails that robust watermarks reside in the global spectral amplitudes. We now analyze how the different schemes achieve this while ensuring stealthiness:
@@ 4.1.2. Spectral Watermarks
Figure 2. Images watermarked by StableSignaturenon-semantic (top), and StegaStampsemantic (bottom). The rightmost figures display the differences between the original and watermarked images that correspond to the changes encoding the watermarks. StableSignature’s modifications are restricted to existing (high-frequency) edges such as wrinkles, hair, mustache, and intersections of multiple components. StegaStamp’s watermark is distributed across the image. The magnified area shows how it manipulates the consistency (texture), injecting gradual (low-frequency) changes that manifest as wrinkles at this location.
@@
Non-Semantic: As they cannot change the content, these are typically injected in high frequencies since sufficiently manipulating lower bands s.t. the watermark is detected requires substantial changes due to the vast energy concentration in this range. To alter the collective distribution of high frequencies, polarization at existing edges can be controlled by adjusting the intensities of the various intersecting components to amplify or attenuate the transitions to their environments. This significantly impacts higher frequency amplitudes via small (invisible) modifications only due to the naturally low energy distribution in this range and by exploiting the characteristics of these regions (edges) that already exhibit high-frequency shifts to mask the watermark’s footprint. Fig. 2 (top) depicts this behavior for the non-semantic StableSignature (Fernandez et al., 2023). We verified that other non-semantic schemes operate similarly.
@@
Semantic: Since they can influence the content, they will populate the consistency-defining low frequencies, making them more robust. As spectral amplitudes represent the collective contribution to pixel variations occurring at the relevant rates from all spatial locations, these schemes construct their outputs s.t. the needed changes are spread among different regions, reducing visual cues. Unlike non-semantic watermarks, they are not restricted to edges, affecting various objects’ structures and textures. Since the changes required to alter low-frequency magnitudes are large, their influence may be visible. Fig. 2 (bottom) portrays this for StegaStamp (Tancik et al., 2020). TRW’s (Wen et al., 2023) images may display different content from the original (non-watermarked), making illustrations uninformative (see §2.1). Yet, our results in §6.2.2 prove its watermarks are in low frequencies.
5. UnMarker
@@
Here, we present our universal attack. It follows from §4.1.1 that the carrier must be a set $\mathcal{\boldsymbol{F}}_{c}=\{|F_{b}|:\;b\in[\![B]\!]\}$ of spectral amplitudes and the watermark property distribution, denoted as ${\mathcal{\boldsymbol{F}}}_{w}$, is over these amplitudes. Let $\ell_{d}$ be a perceptual distance metric s.t. if $\ell_{d}(x,\;y)\leq t_{\ell_{d}}$ for some threshold $t_{\ell_{d}}$, then $x$ and $y$ are visually similar. We can now formalize watermark removal:
@@
(1) $\begin{split}x_{nw}=\underset{x^{\prime}}{argmax}&\;[\ell_{F}(x_{w},\;x^{\prime})]\\[-5.0pt] &s.t.\;\ell_{d}(x_{w},\;x^{\prime})\leq t_{\ell_{d}}\end{split}$
@@
5.1. Attacking Non-Semantic Watermarks
@@
Spectral Loss $\ell_{F}$. We propose the Direct Fourier Loss (DFL):
@@
Norm Constraint: To further ensure no such peaks appear (especially for low resolutions), we also use the $\|\|_{2}$ norm to place a geometric bound on the distance from the watermarked image, allowing us to slightly increase the budget used for LPIPS. Whenever LPIPS risks introducing such peaks, $\|\|_{2}$ grows, forcing the procedure to explore different directions. Yet, $\|\|_{2}$ alone is insufficient since geometric constraints cannot capture visual similarity and will either lead to sub-optimal solutions if used with extremely small thresholds or impact quality otherwise. This is especially true since we operate in a black-box framework without model feedback, which raises the bar and requires larger perturbations, entailing that the perceptual loss should be accurate to preserve quality.

### v2

@@ 3 Threat Model
Providers’ Goals. GenAI providers must ensure their images are not used to spread disinformation. Thus, they watermark generated images. For a suspicious image, they invoke the detector to test for the watermark, proving it fake.
@@
Providers’ Capabilities. They have resources and data to train SOTA generators. As is common practice [1, 3], their APIs are restrictive, accepting queries (e.g., text) and returning watermarked images, hiding intermediate results.
@@
Attacker’s Goals. Attackers wish to generate real-looking deepfakes by invoking the providers’ APIs and then using methods to 1) remove watermarks while 2) retaining quality.
@@
Attacker’s Capabilities: Resources. Following the defensive watermarking assumptions (see §2), attackers lack high-quality data or extensive resources. Thus, training advanced models or surrogates for the providers’ generators [32] or detectors [29] is not possible. As some services (e.g., AWS) allow renting advanced hardware, attackers may access powerful machines. Yet, this can only be short-term due to high costs, not for training large models over long GPU hours.
@@
Black-Box. As many providers keep their algorithms and datasets secret [49], attackers have no whitebox access to (similar) models, data, or any information about the system.
Query-Free. When providers restrict access to their detector to trusted corporations or employees, as confirmed by recent announcements [50], attacks relying on detector feedback [30] lose their efficacy. Given these announcements and defensive watermarking’s vulnerability when query access to the detector is publicly available (see §2.2), we follow previous work [33, 35, 29] and investigate this paradigm against query-free attacks where attackers cannot interact with the detector directly to observe its decisions.
4 Design Philosophy
@@
Since the detector $\mathcal{D}$ must be able to verify the watermark in any image, it should know where to find this watermark, meaning there must be a set of attributes, namely the universal carrier, that can be measured in all images to make this decision. Our analysis will reveal that identifying this set is possible since the requirements from a robust scheme eliminate all options except the spectral amplitudes. Thus, any robust scheme will inject its watermarks in the collective magnitudes of the different frequency bands. In §5, we show how identifying this carrier enables attackers to mount destructive attacks that remove the watermark with minimal visual impact and no knowledge of the system.
4.1 Uncovering The Universal Carrier
@@
Any computationally limited detector $\mathcal{D}$ can only monitor a finite set of attributes $\mathcal{\boldsymbol{P}}=\{p_{b}|\;b\in[\![B]\!]\}$ in all images. Deciding if an image is watermarked relies on whether (a subset of) $\mathcal{\boldsymbol{P}}$’s attributes follow a watermark property distribution $\mathcal{\boldsymbol{P}}_{w}$ that $\mathcal{D}$ associates with the watermark $w$. While $\mathcal{\boldsymbol{P}}$ is a set of attributes, $\mathcal{\boldsymbol{P}}_{w}$ is a distribution over $\mathcal{\boldsymbol{P}}$’s values, that identifies $w$. There may be several patterns $\mathcal{\boldsymbol{P}}_{w}$ pertaining to subsets of $\mathcal{\boldsymbol{P}}$. We term $\mathcal{\boldsymbol{P}}$ the universal carrier as it contains all the properties $w$ can influence. Although $\mathcal{\boldsymbol{P}}_{w}$ should allow watermarking different contents, its variance must still be sufficiently low to prevent detecting $w$ in non-watermarked images, restricting it to a reasonably narrow vicinity. Thus, introducing small deviations into $\mathcal{\boldsymbol{P}}$’s values can cause any watermarked image to violate this distribution $\mathcal{\boldsymbol{P}}_{w}$, removing the watermark without degrading quality. This destructive strategy requires no knowledge of $\mathcal{\boldsymbol{P}}_{w}$, making it feasible without feedback from $\mathcal{D}$ if the carrier $\mathcal{\boldsymbol{P}}$ is correctly identified. To identify $\mathcal{\boldsymbol{P}}$, we analyze widely-accepted requirements [18, 23, 19, 24, 21, 22, 20] from a robust scheme:
@@
Robust Watermarking Requirements. Universal Applicabil- ity: The scheme must watermark images of various concepts and scenes, and the watermark must always be detectable.
@@
Universal Stealthiness. The process cannot introduce unnatural artifacts that degrade quality. We do not require the watermark not to leave visible traces since semantic watermarks violate this concept by definition, influencing the structure. Yet, as users only observe watermarked images, quality becomes the only relevant metric for stealthiness.
Figure 1: Spectral analysis of two images. Both depict worldly objects of high consistency. Thus, the collective spectral magnitudes are distributed similarly as the low frequencies (center) corresponding to gradual variations are always far more dominant. Phases determine the spatial arrangement of the pixel value shifts that constitute these magnitudes to shape the content, making them extremely different.
@@ 4.1.1 Candidates
An image is characterized by its 1) spatial structure (pixel values), and 2) frequency distribution— the rates at which values alternate between neighboring pixels. Thus, the carrier must reside in these channels. Below, we present hypotheses, accompanied by an in-depth analysis, regarding the possibility of injecting watermarks in each of these channels, identifying a unique carrier capable of satisfying all demands. Our claims are corroborated by UnMarker’s empirical success in §6.
@@
H1. Watermarks cannot be embedded in the spatial structure (pixel values). This is because they make for an unreliable carrier that experiences sporadic variations (in the absence of a watermark) between images based on their contents (different images assign extremely different values to their pixels). However, the scheme must always bring the carrier’s values to the same distribution $\mathcal{\boldsymbol{P}}_{w}$ for any watermarked image. When these values naturally (i.e., without the watermark) exhibit a high degree of content dependency, this is impossible without restricting the scheme to processing a small set of similar contents or leaving unnatural cues in many outputs corresponding to contents that would otherwise require these values in images portraying them to lie far away from the desired $\mathcal{\boldsymbol{P}}_{w}$. As this distance increases, the magnitudes of the changes needed to bring them to $\mathcal{\boldsymbol{P}}_{w}$ grow, failing to integrate naturally with the content. Restrictions violate applicability, while unnatural artifacts are not stealthy. The scheme may strategically place pixels encoding the watermark to accommodate large changes with minimal visibility. Two possibilities exist: either they are far apart to remain unnoticeable despite significantly differing from their environments or located in regions the content is unlikely to occupy (i.e., the background). Both are volatile and cannot withstand manipulations (e.g., cropping or noise), violating the robustness requirement (see §6.2.1).
@@
Figure 2: Images watermarked by StableSignaturenon-semantic (top), and StegaStampsemantic (bottom). The rightmost figures display the differences between the original and watermarked images that correspond to the changes encoding the watermarks. StableSignature’s modifications are restricted to existing (high-frequency) edges such as wrinkles, hair, mustache, and intersections of multiple components. StegaStamp’s watermark is distributed across the image. The magnified area shows how it manipulates the consistency (texture), injecting gradual (low-frequency) changes that manifest as wrinkles at this location.
We now turn our focus to the frequency distribution. There are two options for influencing frequencies to inject watermarks: Locally or globally. We claim the following:
H2. Local changes to the pixel shift rates cannot reliably encode watermarks. The reason is that local modifications affect the rates at which pixels vary in select spatial regions. These share the above pitfall of strong dependency on the content, leaving only global spectral changes. The spectrum associates frequencies with two properties: magnitude and phase. Magnitudes represent the prevalence of shifts that occur at the specific rate across the image. Mathematically, any image is a superposition of periodic signals with varying frequencies occurring at different phases that determine their orientations, forming the image’s structure based on how they intersect. Accordingly, and similar to the spatial structure, we deduce that: H3. Phases are an unreliable carrier due to their high sensitivity to the content. This is the result of the phases varying drastically between images as the primary factor affecting the structure (see Fig. 1).
By process of elimination, hypotheses H1-H3 exclude all other candidates, leaving only one potential carrier: the global spectral magnitudes. These are collective measurements with little dependency on specific pixels, making them resistant to reasonable manipulations that would preserve the majority of energy (amplitudes) distribution. They are also distributed similarly for any realistic content: lower frequencies contain the majority of energy as they correspond to gradual changes that capture the inherent consistency (continuity) of the worldly objects images depict. Higher frequencies (rapid transitions) occur at edges (intersections of objects), making them far less prevalent (see Fig. 1). Thus, while amplitudes vary between images, they always follow a similar distribution, decaying as the frequencies grow. This stability other alternatives lack allows injecting watermarks via changes of similar magnitudes for any content, enabling them to be minimized s.t. they always integrate naturally.
4.1.2 Spectral Watermarks
The above entails that robust watermarks reside in the global spectral amplitudes. Based on the different schemes’ characteristics, we now present our main hypothesis, which serves as UnMarker’s foundation:
H4. Non-semantic schemes primarily inject watermarks in the global spectral amplitudes associated with high frequencies, while semantic watermarks mainly target low frequencies. Below, we analyze this claim:
@@ 4.1.2 Spectral Watermarks
Non-Semantic: As these schemes cannot change the content, these are typically injected in high frequencies since sufficiently manipulating lower bands s.t. the watermark is detected requires substantial changes due to the vast energy concentration in this range. To alter the collective distribution of high frequencies, polarization at existing edges can be controlled by adjusting the intensities of the various intersecting components to amplify or attenuate the transitions to their environments. This significantly impacts higher frequency amplitudes via small (invisible) modifications only due to the naturally low energy distribution in this range and by exploiting the characteristics of these regions (edges) that already exhibit high-frequency shifts to mask the watermark’s footprint. Fig. 2 (top) depicts this behavior for the non-semantic StableSignature [20]. We verified that other non-semantic schemes operate similarly.
@@
Semantic: Since these schemes can influence the content, they will populate the consistency-defining low frequencies, making them more robust. As spectral amplitudes represent the collective contribution to pixel variations occurring at the relevant rates from all spatial locations, these schemes construct their outputs s.t. the needed changes are spread among different regions, reducing visual cues. Unlike non-semantic watermarks, they are not restricted to edges, affecting various objects’ structures and textures. Since the changes required to alter low-frequency magnitudes are large, their influence may be visible. Fig. 2 (bottom) portrays this for StegaStamp [24]. TRW’s [18] images may display different content from the original (non-watermarked), making illustrations uninformative (see §2.1). Yet, our results in §6.2.2 prove its watermarks are in low frequencies.
5 UnMarker
@@
Now that we have identified the carrier, we present our universal attack. It follows from §4.1.1 (H4) that the carrier must be a set $\mathcal{\boldsymbol{F}}_{c}=\{|F_{b}|:\;b\in[\![B]\!]\}$ of spectral amplitudes and the watermark property distribution, denoted as ${\mathcal{\boldsymbol{F}}}_{w}$, is over these amplitudes. Let $\ell_{d}$ be a perceptual distance metric s.t. if $\ell_{d}(x,\;y)\leq t_{\ell_{d}}$ for some threshold $t_{\ell_{d}}$, then $x$ and $y$ are visually similar. We can now formalize watermark removal:
@@
[…]
@@
5.1 Attacking Non-Semantic Watermarks
@@
Spectral Loss. We propose the Direct Fourier Loss (DFL):
@@
Norm: To further ensure no such peaks appear (especially for low resolutions), we use the $\|\|_{2}$ norm to place a geometric bound on the distance from the watermarked image, allowing us to slightly increase the budget used for LPIPS. Whenever LPIPS risks introducing such peaks, $\|\|_{2}$ grows, forcing the procedure to explore different directions. Yet, $\|\|_{2}$ alone is insufficient since geometric constraints cannot capture visual similarity and will either lead to sub-optimal solutions if used with extremely small thresholds or impact quality otherwise. This is especially true since we operate in a black-box framework without model feedback, potentially requiring larger perturbations, entailing that the perceptual loss should be accurate to preserve quality.